OpenStudy (anonymous):

How can I decide whether this integral is convergent or divergent? ∫(0,1) 1/(x^3-2x^2+x)

OpenStudy (anonymous):

I have a phrase they say if: The integral ∫ (0.1) 1/x^p converges if p <1 and diverges if p ≥ 1 I'm almost sure it's the one I should use, but I do not know how.

OpenStudy (john_es):

I would try to "factorize", \[\int_0^1\frac{1}{x^3-2x^2+x}dx=\int_0^1{\left(\frac{1}{x}-\frac{1}{x-1}+\frac{1}{(x-1)^2}\right)dx}\] And apply the theorem term by term,

OpenStudy (anonymous):

Thank you. If each term diverges so does the integral it too? You know what to do if some of the terms converge and other diverge?

OpenStudy (agent0smith):

If any one term diverges, the whole integral does.

OpenStudy (agent0smith):

"∫ (0.1) 1/x^p converges if p <1 and diverges if p ≥ 1" Isn't this the other way around? It diverges if p =< 1 and converges if p>1.

OpenStudy (anonymous):

I dont think so... Wacth the attachment.

OpenStudy (anonymous):

Agent0smith Perhaps you are thinking of the integral: \[\int\limits_{1}^{∞}dx/x^p\] converges if p > 1 and diverges if p ≤ 1. ???

OpenStudy (agent0smith):

Yeah, it appears you're right. http://www.sosmath.com/calculus/improper/convdiv/convdiv.html

OpenStudy (john_es):

If one of the terms diverges, then the integral diverges. In fact, you can see that the last term is the one that make the true divergence, numerically spaking. An alternative way to test the divergence is calculating the integral. Then you will see that each term diverges. If one term converges but one diverges, then the integral diverges.

OpenStudy (anonymous):

John have won solve this integral: \[\int\limits_{0}^{1}1/(x^3+x^2-2x)\]?

OpenStudy (anonymous):

maybe someone else can help me?

OpenStudy (agent0smith):

\[\Large \int\limits_0^1{\left(\frac{1}{x}-\frac{1}{x-1}+\frac{1}{(x-1)^2}\right)dx}\] All terms of this are divergent.

OpenStudy (anonymous):

Thank you Smith

OpenStudy (anonymous):

I do not think the equation is correct. It's the same as previously!?!

OpenStudy (john_es):

The indefinite integral can be calculated, \[\ln|x|-\ln|x-1|-\frac{1}{x-1}+K\] You can check the limits x->0, and x->1, you will find the expression is divergent.

OpenStudy (agent0smith):

@RFJ-86 the two are equal, you can check it by getting the left side over a common denominator x(x-1)^2 (John_ES used partial fractions to break it up) \[\LARGE {\frac{1}{x}-\frac{1}{x-1}+\frac{1}{(x-1)^2}} = \frac{1}{x^3-2x^2+x}\]

OpenStudy (agent0smith):

\[\Large \int\limits\limits_0^1{\left(\frac{1}{x}-\frac{1}{x-1}+\frac{1}{(x-1)^2}\right)dx} \] From this all you need to realize is that 1/x is divergent, so the whole integral is (if one term diverges to infinity, does it matter if the other terms converge?)

OpenStudy (anonymous):

@agent0smith I agree with this equation \[\frac{ 1 }{ x }-\frac{ 1 }{ x-1 }+\frac{ 1 }{ (x-1)^2 }=\frac{ 1 }{ x^3-2x^2+x }\] But this is one i not true: \[\frac{ 1 }{ x }-\frac{ 1 }{ x-1 }+\frac{ 1 }{ (x-1)^2 }\ \neq \frac{ 1 }{ x^3+x^2-2x }\] I need a way to "factorize" this equation \[\frac{ 1 }{ x^3+x^2-2x }\]

OpenStudy (agent0smith):

But your original integral was the 1/(x^3-2x^2+x)... For the one you just posted, pull out an x \[\Large \frac{ 1 }{ x(x^2+x-2) }\] then factor the brackets \[\Large \frac{ 1 }{ x(x+2)(x-1) }\] Now you should be able to make partial fractions.

OpenStudy (anonymous):

Thank you @agent0smith